What's the rate of change, y intercept and x intercept for these equations y=6-2/3x, 3(2x-8)=6y, 120=y+4x
Answer by stanbon(75887)

What's the rate of change, y intercept and x intercept for these equations
y=6-2/3x,
If x = 0, y = 6
If y = 0, x = (3/2)6 = 9
rate of change or slope = (-2/3); y-intercept = 6 ; x-int = 9
-----------------------------
3(2x-8)=6y,
y = (1/2)(2x-8) = x-4
If x = 0, y = -4
If y = 0, x = 4
slope = 1, y-int = -4, x-int = 4
-----------------------------
120=y+4x
y = -4x + 120
If x = 0, y = 120
If y = 0, x = 30
slope = -4, y-int = 120, x-int = 30
-----------------
Cheers,
Stan H.
-----------